70 coins of 10 paise and 50 paise are kept in a purse. If yhe total value of the money in the purse is rs 19. Find the number of each types coin.
Answer by FrankM(1040)

you have 70 coins of 10 and 50 value, total value is 1900.
I will call the 10 unit T, and 50 unit, F, in other words, T is the number of 10 units coins.
T+F=70
10T+50F=1900
Now multiply first equation by -10 , both sides,
-10T-10F=-700
10T+50F=1900 and add the second equation
40F=1200
F=30
T=40
40*10 =400
30*50=1500
Total value 1900